# Guided Decoding for Robot Motion Generation and Adaption

### v1

@@ IV-D Adaptation and generalization
The experiments are trained on a Panda robot. Our adaption and generalization are suitable for both samples from priors and posterior. To enhance visualization and compare with demonstrations, we present all results based on samples from posteriors, except for those in Sections IV-D2 and IV-D4 which also use samples from priors.
IV-D1 Via points
IV-D2 Obstacle avoidance
IV-D3 Position bound
IV-D4 Velocity and acceleration bound
@@
We have established boundaries for both velocity and acceleration, as detailed in Table I. Although the table presents only a single boundary for each type of boundary, these boundaries can be set to arbitrary values.

### v2

@@ IV-D Adaptation and generalization
The experiments are trained on a Panda robot. Our adaption and generalization are suitable for both samples from priors and posterior. To enhance visualization and compare with demonstrations, we present all results based on samples from posteriors, except for those in Sections IV-D4 and IV-D2 which also use samples from priors.
IV-D1 Position bound
IV-D2 Velocity and acceleration bound
We have established boundaries for both velocity and acceleration, as detailed in Table LABEL:tab:v_a_limit. Although the table presents only a single boundary for each type of boundary, these boundaries can be set to arbitrary values.
IV-D3 Via points
IV-D4 Obstacle avoidance